A novice skier, starting from rest, slides down a frictionless 35 degree incline whose vertical height is 178 m. How fast (m/s) is she going when she reaches the bottom?


1
Expert's answer
2021-03-02T18:04:02-0500
"KE_i+PE_i=KE_f+PE_f,""0+PE_i=KE_f+0,""mgh=\dfrac{1}{2}mv_f^2,""v_f=\sqrt{2gh}=\sqrt{2\cdot 9.8\ \dfrac{m}{s^2}\cdot178\ m}=59.1\ \dfrac{m}{s}."